OpenStudy (anonymous):

HELP PLEASE !!!!!!!!!!!!!! Solve for x by completing the square: a. x^2-mx-6m^2=0 b. kx^2+px+q=0

OpenStudy (anonymous):

\[a.\ x ^{2}-mx-6m ^{2}=0\] \[b. \ kx^{2}+px+q=0\]

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

a) x^2-mx=6m^2....now add the square of (m/2) to both sides x^2-mx+m^2/4=6m^2+m^2/4 (x-m/2)^2=25m^2/4 can you take it from here?

OpenStudy (anonymous):

Guadratic formula: \(\huge \frac{ - \color{blue}{b} \pm \sqrt{\color{blue}{b^2} -\color{black}{4} \color{green}{a} \color{red}{c} }}{\color{black}{2} \color{green}{a}}\) For Guadratic Equation \(\Large \color{green}{a} \color{black}{x^2}+\color{blue}{b} \color{black}{x}+\color{red}{c}=\color{black}{0}\) 1) \(\Large x^2-mx-6m^2=0\) \(\huge \Rightarrow x=\frac{m\pm \sqrt{(-m)^2-4\times (-6)}}{2}=\frac{m\pm \sqrt{m^2+24}}{2}\) 2) \(\Large kx^2+px+q=0 \) \(\huge \Rightarrow x=\frac{-p\pm \sqrt{p^2 -4kq}}{2k}\)

OpenStudy (anonymous):

Can somebody translate it like for a year 9 student please? :-) thanks you

OpenStudy (anonymous):

use mine its g9 friendly